5. Aluminium (Group III) and Silicon Group (IV) are adjacent to each other on the periodic table. Relative to aluminium, silicon is less dense and has a higher melting temperatures, is harder and is very prevalent in sand and rocks. Yet, aluminium has more mechanical applications in engineering systems while Si finds more applications in electronic systems Explain how the functional properties of these element arise and why they are suited for different functions. 6. Suppose an element has a valence of 2 and an atomic number of 27. Based only on quantum numbers, how many electrons must be present in the 3d energy level? 7. Compare and contrast metallic and covalent primary bonds in terms of: (a) the nature of the bond (b) the valence of atoms involved (c) electrical conductivity of solids bonded this way (d) the ductility of materials bonded this way 8. Calculate the fraction of bonding of MgO that is ionic. 9. Which of MgO or magnesium would expect to have higher melting temperature? Explain. 10. Aluminium and Silicon are side by side on the periodic table. Which of the two would you 11. Steel is coated with thin layer of ceramic to help protect against corrosion. What do you expect to have higher modulus of elasticity? Explain expect to happen to the coating when the temperature of the steel is increased significantly? Explain.
Answer #1

5. Aluminium is a metal whereas Silicon is a metalloid.

Aluminium alloys are used in structures and components where light weight or corrosion resistance is required.

Silicon on the other hand being a metalloid has the ability to be semiconductor because of 14 electrons silicon need four electrons to become stable so it undergoes covalent bonding. Silicon can be doped with either electron rich compound containg 15 electrons or electron deficient compound containing 13 electrons to form a n- type and p-type semiconductors respectively. Semiconductors have very vast application in electrical systems like transistors, etc.

6. For electronic configuration we follow n+l rule which means that the orbital having larger value for n+l will be higher in energy and electrons are filled in orbitals following energy order. By this rule 3d has n+l = 3+2=5 whereas 4s has n+l = 4+0 = 4.

So 4s will be filled before 3d

electronic configuration will be,

1s2 2s2 2p6 3s2 3p6 4s2 3d7 so ,

The number of electrons present in 3d energy level are 7

Please cilck on thumbs up if this answer is of any help.